# Scalable hybrid quantum Monte Carlo simulation of U(1) gauge field coupled to fermions on GPU

### v3

@@ I Results
Figure 3: The correlation functions for non-compact QED3 with $K=1$ and $J=1.25$. (a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$, for various cubic lattice sizes shown as $N_{\tau}\times L^{2}$ in the legend. Notice that $C_{S}^{\uparrow\downarrow}=C_{S}^{\downarrow\uparrow}$, so we only need to measure $C_{S}^{\uparrow\downarrow}$ here. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray square, triangular and diamond dots show the results computed from the DQMC to benchmark the HQMC result of the corresponding lattice sizes. At the large distance and the large system sizes, the data are manually fitted by the black solid line, which displays a power law decay as $1/r^{3.8}$. (b) shows the log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$. The legend is the same as that in (a). At the large distance and the large system sizes, the manually fit line of the bond-bond correlation also displays a power law decay as $1/r^{3.7}$. (c) shows the log plot of the flux-flux correlation as a function of imaginary time $\tau$. At the large distance and the large system sizes, the black solid line shows the manual fitting which displays a power law decay as $\tau^{-4.0}$. The black dashed line shows the $\tau^{-3.0}$ power law behavior for comparison.
@@
Fig. 3(a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of real space distance $r$ at $J=1.25,K=1$, for various lattice sizes. As the system size increases from $L=12$ all the way to $L=60$, the spin correlation curves monotonically unroll, and approach a straight envelope line at large distances. This indicates a power law decay behavior. By manually fitting this envelop curve shown as the dark solid line, we obtain the exponent of the power law decay to be around $3.8$. This agrees with the scaling dimension of the fermion bilinear operators $2\Delta_{\textmd{adj}}\in(3.5,3.9)$ predicted by the CFT, where the fermion bilinear operators serve as the adjoint representation of the emergent $SU(4)$ symmetry. We note that at around $1\times 10^{-5}$ in Fig. 3(a), the data fall below the noise level. So the manual fitting is performed mainly with the few large-distance data $\gtrsim 10^{-5}$ at large system sizes. In Fig. 3(a), we also plot the DQMC results for several small lattice sizes to benchmark the HQMC results, and they well agree with each other. These results are complemented by the benchmark data in Sec.II of the Supplemental Materials (SM) [suppl].
@@
Fig. 3(b) shows log plot of the bond-bond correlation $C_{B}(r,0)$ as a function of real space distance $r$ at $J=1.25,K=1$, for various lattice sizes. The data of the bond correlation generally have a larger variance than the spin correlation function shown in Fig. 3(a), which is due to the explicit gauge field dependence in the bond operator $B_{i}$. But as the system sizes increases to beyond $L\gtrsim 30$, the decaying behavior of the bond-bond correlation tends to be stable around a straight line, which again indicates a power law decay. The manual linear fitting shown as the dark solid line yields a power of around $3.7$, which again agrees with the scaling dimension of the fermion bilinear operators $2\Delta_{\textmd{adj}}\in(3.5,3.9)$.
@@
As analyzed in the above section and shown in Figs. 3(a) and 3(b), the fact that both spin and bond correlation functions display power law decay with powers consistent with the scaling dimension of the CFT fermion bilinear operators supports the existence of the $U(1)$ DSL phase in non-compact QED3 at $J=1.25$. This also agrees with a recent DQMC study on the same model [chenEmergent2025].
@@
Fig. 3(c) shows the log plot of the magnetic flux correlation function $C_{\textmd{flux}}(\tau)$ as a function of imaginary time $\tau$ for various lattice sizes, at $J=1.25,K=1$. It can be seen that, as system size increases, the curves become increasingly linear before entering the noise level at around $2\times 10^{-3}$. Such linear tendency is manually fitted with the dark solid line, which yields a power of around $4.0$. In contrast, the black dashed line shows the $\tau^{-3.0}$ behavior, which originates from the non-conformal flux-flux correlation in a $U(1)$ gauge field theory [chenEmergent2025]. This is clearly distinct from the $\tau^{-4.0}$ behavior originating from the conserved current’s scaling dimension $\Delta_{J}=2$, as mentioned above. Thus, the power law decay in the flux correlation function with power $4.0$ also indicates the existence of $U(1)$ DSL phase in $J=1.25$ in the non-compact QED3.
@@ II Discussion
Since the HQMC gives the full space-time correlation of the fermion bilinear operators, one immediate future direction is to investigate the spectral information inside the $U(1)$ DSL phase from the stochastic analytic continuation of the HQMC data [chenEmergent2025]. As we have both large $N_{\tau}$ and $V_{s}$, the momentum and energy resolution of the obtained spectra (for example the magnetic spectra from spin-spin correlation function) will be much higher than those from the DQMC [chenEmergent2025]. In this way, the weight distribution and the momentum dependence of the spinon continuum spectra of $U(1)$ DSL can be revealed, and hopefully connection with experiment results on triangular lattice antiferromagnets YbZn2GaO5 [BagEvidence24, WuSpinDynamics25] and the $A$-YbSe2 delafossites [ScheieKYbSe2, ScheieNaYbSe2] and kagomé antiferromagnet YCu3(OD)6Br2[Br1-x(OD)x[zengPossible2022, zengSpectral2024, hanSpin2024, suetsuguEmergent2024, suetsuguGapless2024, jeonOne2024, zhengUnconventional2025] can be made.
Moreover, by adding magnetic field to the QED3 model, one can investigate how the flavor chemical potential will change the $U(1)$ DSL phase and understand the emergent gauge flux in magnetized Dirac spin liquids. For example, recent DQMC work [chenEmergent2025] and analytic works [dumitrescuSymmetry2024, dumitrescuComments2025] have explored the various symmetry-breaking phases out of $U(1)$ DSL under magnetic field. In particular, a novel chiral flux phase, where the system spontaneously generates a net emergent orbital magnetic flux (deviated from the $\pi$-flux at zero field case) has been seen from smaller size DQMC simulation. This spontaneous orbital magnetic flux changes the behavior of the spinons, and make them form the relativistic Landau levels — energy levels normally seen in electrons moving in a magnetic field. Remarkably, the system also exhibits a gapless photon-like mode, a signature of the coherent emergent gauge field, which can be potentially detected in inelastic neutron scattering experiments [chenEmergent2025]. Our GPU-accelerated HQMC algorithm can provide systematic results of such interesting phase and rich diagram, as a function of magnetic field, gauge field fluctuations, as well as the temperature axes. One possible direction is to investigate the effect of different filling of the spinon Landau levels in the magnetized DSL phase [SodemannPseudoscalar21], and establish possible connections to experimentally observed unusual magnetic oscillations in the YCu3(OD)6Br2[Br1-x(OD)x] kagomé antiferromagnet [zhengUnconventional2025].
@@ III Methods
As introduced above in the Results section, the model we study consists of a fermionic Lagrangian $L_{F}$, and a gauge field Lagrangian $L_{B}$. And we consider two $U(1)$ gauge field Lagrangians, namely compact $U(1)$ gauge field $L_{B}^{\textmd{c}}$ and non-compact $U(1)$ gauge field $L_{B}^{\textmd{nc}}$ as introduced in the Results section. The discretized version of $\operatorname{curl}\phi$ for a spatial plaquette $\square$[degrand1980topological, Fiebig1990] , which is also seen as magnetic flux through the plaquette, is defined as $\theta^{xy}_{\square}=\operatorname{curl}\phi=\sum_{ij\in\textmd{edge}(\square)}\phi_{ij}$, where $ij$ is along the positive direction, and $\phi_{ij}=-\phi_{ji}$, as illustrated in the bottom right cube in Fig. 1. Similarly, in the cubic space-time lattice, the curl of a temporal plaquette, i.e. $\theta^{\tau x}$ or $\theta^{\tau y}$, can also be defined. If we fix the gauge field of temporal bonds to be zero, then the curl of a temporal plaquette is $\pm(\phi_{ij,\tau+1}-\phi_{ij,\tau})$, which consists of the $\tau$-derivative term in the gauge field Lagrangians. It is then clear that $L_{B}^{\textmd{c}}$ has $2\pi$-periodicity in the magnetic flux: $\theta\to\theta+2n\pi,n\in\mathbb{Z}$. Thus, $L_{B}^{\textmd{c}}$ describes a compact $U(1)$ gauge field theory in the sense that the magnetic flux is physical only in the range of $(-\pi,\pi]$, while the other values of $\theta$ are unphysical and need to be wound into the $(-\pi,\pi]$ interval. Due to $2\pi$-periodicity (the fermionic Lagrangian $L_{F}$ also has $2\pi$-periodicity by definition [lieb1994flux]), the compact QED3 is known to host magnetic monopoles, which contains distinct physics from a non-compact theory e.g. $L^{\textmd{nc}}_{B}$ in Eq. (2), where the first term breaks the $2\pi$-periodicity. For example, the proliferation of magnetic monopoles in compact QED3 will lead to the emergence of a mass gap in the gauge field, through the dual superconductor mechanism [polyakov1975compact, polyakov1977quark, karthikNumerical2019], while the non-compact QED3 is generally believed to remain deconfined. We therefore focus on the results on non-compact QED3 in the main text and leave the compact QED3 results in the Supplemental Materials (SM) [suppl]. Although from the algorithmic point of view, there is no difference in their implementation and performance of the corresponding simulations.
@@
is of size $N_{\tau}V_{s}\times N_{\tau}V_{s}$. Furthermore, it can be shown [XYXuU12018] that $\operatorname{det}M_{\alpha}\in\mathbb{R}$ and $\operatorname{det}M_{\uparrow}=\operatorname{det}M_{\downarrow}$, thus the simulation is free from the sign problem. Based on the formula above, the partition function can then be expressed as
@@
In this subsection, we will briefly introduce the HQMC algorithm applied to simulate the partition function Eq. (9). This algorithm originates from lattice quantum chromodynamics simulation [Duane1987], and later finds its application in a broader class of statistical problems [brooks2011handbook]. An application in the Hubbard model setting of condensed matter physics can be found in Ref. [Beyl2018]. In the most recent progress, the method is significantly accelerated with the help of GPU in the study of spin-fermion models and the associated non-Fermi-liquid and quantum critical metallic states [patel2024strange, lunts2023non].
@@
Here, the Gaussian integral of the complex fields $\eta\in\mathbb{C}^{V_{s}N_{\tau}}$ has been conversely used, and the identity $\operatorname{det}M_{\uparrow}=\operatorname{det}M_{\downarrow}\in\mathbb{R}$ is assumed, and proved in Ref. [XYXuU12018]. Thus, the fermionic flavor $\alpha$ is omitted.
@@
First, the Hamiltonian dynamics has the property that the phase space volume is conserved, according to the Liouville theorem [Beyl2018], which ensures that $H(\phi_{0},p_{0},\eta)=H\left(\phi_{t},p_{t},\eta\right)$ and that the proposal variables $(\phi_{t},p_{t})$ are almost always accepted, if the differential equation is solved exactly.
Second, we apply the well-known Leapfrog method [Beyl2018] to solve the differential equation Eq. (14). It has the property of time reversibility, which, together with the Metropolis-Hasting rule, guarantees that the Monte Carlo updates fulfill the detailed balance condition [Beyl2018]. The Leapfrog method is described as follows. At the $n$-th step, to evolve a given $(\phi_{n-1},p_{n-1})$ for a time step $\Delta t$,
@@
Third, since the update scheme in HQMC is based on solving the Hamiltonian differential equation, it is naturally continuous. As a result, the random walk in HQMC is limited to a continuous region the configuration space. But in the QED3 model studied in this paper, there are two non-continuous regions, namely those with $(\det M_{\uparrow},\det M_{\downarrow})=(+,+)$ and $(\det M_{\uparrow},\det M_{\downarrow})=(-,-)$. They are separated by the gauge configurations where $\det M\approx 0$, which themselves have diminishing contribution in the partition function Eq. (9), but prevents the HQMC update from going from one region to the other. This leads to the ergodicity issue, shared by the application of HQMC on other models [Beyl2018] as well. One common solution is to introduce another legitimate update that is ergodic. Then the combination of the multiple updates will cure the ergodicity issue in HQMC [brooks2011handbook]. In Ref. [Beyl2018], the solution of this issue is to decouple the Hamiltonian into two channels and then to complexify the configuration space, which is more problem-specific. In the QED3 studied here, we find that in the small $J$ region, which is the $U(1)$ DSL phase region and the focus of this paper, the ergodicity issue is numerically negligible. So we directly apply the HQMC algorithm in this region. The validity is benchmarked by the DQMC algorithm and is shown in Sec. II of SM [suppl].
@@
For a given bosonic configuration $\phi$, the fermionic observables are easily computed by using the standard stochastic estimation procedure [Duane1987, cohen2022fast]. More specifically, the single-particle Green’s function, $G_{ij}=\langle c_{i}(\tau)c^{\dagger}_{j}(0)\rangle$, which is obtained from the inversion of the $M$ matrix, $M^{-1}$. The direct inversion of $M$ is numerically expensive and scales as $O((V_{s}N_{\tau})^{3})$. Alternatively, the standard stochastic procedure to compute $M^{-1}$, as an iterative method, is usually more efficient. In this procedure, first sample a random vector $\xi\in\mathbb{C}^{V_{s}N_{\tau}}$, whose entries are i.i.d. and have the orthogonality property $\mathbb{E}_{\xi}\xi^{*}_{i}\xi_{j}=\delta_{ij}$. Then, $\sum_{l}\left[M^{-1}\right]_{il}\xi_{l}\xi^{*}_{j}$ will give an unbiased estimation of $\left[M^{-1}\right]_{ij}$, since
@@
Here, the random vector can be drawn from any distribution that satisfies $\mathbb{E}_{\xi}\xi^{*}_{i}\xi_{j}=\delta_{ij}$, including Gaussian distribution, Rademacher distribution, etc [cohen2022fast]. In practice, we compute $\sum_{l}\left[\left(M^{\dagger}M\right)^{-1}\right]_{il}\left[M^{\dagger}\xi\right]_{l}\xi^{*}_{j}$, instead, where $R:=M^{\dagger}\xi$ is numerically cheap to compute and $\left[M^{\dagger}M\right]^{-1}R$ can be computed generically using a pCG algorithm. In this way, the Green’s function can be obtained, and other more complicated correlation functions can be further obtained from Green’s function via Wick decomposition.
@@
has a time complexity of $O(N^{2})$, where $N=V_{s}N_{\tau}$. Here, we follow the Ref. [cohen2022fast], and apply the fast Fourier transformation (FFT) instead of directly eevaluating this average.
@@
where $\mathcal{F}$ is FFT operation. Then the time complexity is reduced to $O(N\log N)$, with $N=V_{s}N_{\tau}$. Moreover, as introduced in Ref. [cohen2022fast], each random vector among $N_{\textmd{rv}}$ random vectors can serve as either $\xi$ or $\xi^{\prime}$ in the example above. Thus, totally $f(N_{\textmd{rv}})=\binom{N_{\textmd{rv}}}{2}$ random vectors can be effectively used in the stochastic estimation. In our benchmark, $N_{\textmd{rv}}=40$ is sufficient to reduce the error of fermionic bilinear’s correlation function to a moderate level.
@@
As introduced above, the key step in reducing the cubic complexity of computing the determinant $\det M$ or the inversion $M^{-1}$ is to convert it to the evaluation of $\left[M^{\dagger}M\right]^{-1}v$, which is equivalent to solving a linear system $OX=v$. Here, $O=M^{\dagger}M$ is a Hermitian positive semi-definite matrix; thus, the equation can be efficiently solved by the iterative preconditioned conjugate gradient (pCG) method [Beyl2018].

### v4

@@ I Results
Figure 3: The correlation functions for non-compact QED3 with $K=1$ and $J=1.25$. Panel (a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of distance $r$, for various cubic lattice sizes shown as a format of $N_{\tau}\times L^{2}$ in the legends in both panel (a) and (b). Notice that $C_{S}^{\uparrow\downarrow}=C_{S}^{\downarrow\uparrow}$, so we only need to measure $C_{S}^{\uparrow\downarrow}$. To avoid the even-odd oscillation in the finite-size data, we only plot the data of the distance $r=\textmd{odd}$ points. The gray data points show the results computed from the DQMC. The inset plots the zoomed in view of the dashed region, which shows that DQMC results agrees with the HQMC result within the error bars. The HQMC data points at the largest lattice size, above the noise level of $10^{-6}$, are manually fit by the black solid line, which displays a power law decay as $1/r^{3.8}$. Panel (b) shows the log plot of the bond-bond correlation (along $\hat{x}$ direction) as a function of distance $r$ for various lattice sizes. The data points of the largest lattice size, before entering the noise level at $10^{-5}$, are manually fit by the black solid line, and also displays a power law decay with $1/r^{3.7}$. Panel (c) shows the log plot of the flux-flux correlation as a function of imaginary time $\tau$. At the large $\tau$ and the large system sizes, the black solid line shows the manual fitting which displays the clear linear behavior indicating a power-law decay as $\tau^{-4.0}$. As a comparison, the black dashed line shows the $\tau^{-3.0}$ power law behavior. The inset plots zoomed in view of the data in the dashed squared region. Panels (d-f) show the operator scaling exponent $\Delta_{S}$, $\Delta_{B}$, $\Delta_{\textmd{flux}}$ for different lattice sizes obtained from the linear fitting of the data points shown in panels (a-c), respectively (the fit lines are not shown). To obtain the exponent shown in panels (d, e), the data points in panels (a, b) above the noise level for each lattice size are linearly fit, and the slope gives the scaling exponent shown as the blue dots in panels (d, e), which are then extrapolated to $1/L\to 0$, leveraging the rational curves $y=y_{0}+\frac{b\cdot x^{d}}{1+c\cdot x}$ (red dashed curves), where $y_{0},b,c,d$ are free parameters. Similarly, the blue dots in panel (f) are the exponents obtained from the linear fitting of the data points in panel (c) for each lattice size. The data points used in the linear fitting are the linear region above the noise level, which are taken from the $C_{\textmd{flux}}$ interval of $10^{-3},2\times 10^{-2}]$. Again, the red dashed curve and the dot show the $1/L\to 0$ extrapolation.
@@
Fig. 3(a) shows the log plot of the spin-spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ as a function of real space distance $r$ at $J=1.25,K=1$, for various lattice sizes. As the system size increases from $L=12$ all the way to $L=60$, the spin correlation curves monotonically unroll, and approach a straight envelope line at large distances. This indicates a power law decay behavior. By manually fitting this envelop curve shown as the dark solid line, we obtain the exponent of the power law decay to be around $3.8$. This agrees with the scaling dimension of the fermion bilinear operators $2\Delta_{\textmd{adj}}\in(3.5,3.9)$ predicted by the CFT, where the fermion bilinear operators serve as the adjoint representation of the emergent $SU(4)$ symmetry. We note that at around $1\times 10^{-5}$ in Fig. 3(a), the data fall below the noise level. So the manual fitting is performed mainly with the data points $\gtrsim 10^{-5}$ at large system sizes. In Fig. 3(a), we also plot the DQMC results for several small lattice sizes to benchmark the HQMC results, and they well agree with each other. These results are complemented by the benchmark data in Sec.II of the Supplemental Materials (SM) [2].
In order to systematically obtain an estimate of the spin operator scaling exponent along with its error, we perform linear fitting on the log-log plot of the spin correlation $C_{S}^{\uparrow\downarrow}(r,0)$ data for each lattice size in Fig. 3(a) (the fit lines are not shown). The data points above the noise level $10^{-6}$ are all used in the linear fitting. Note that the amount of effective $C_{S}^{\uparrow\downarrow}(r,0)$ data points are few, and do not show clear downward curvature like the $C_{\textmd{flux}}(\tau)$ data shown in Fig. 3(c), so the fitting window starts from the first data points ($r=1$), while the difference resulting from different choices of the fitting window is considered as the error of the fitting slopes, shown as the blue dots error bar in Fig. 3(d). The extracted scaling exponents for each lattice size are shown as the blue dots in Fig. 3(d), which are then fitted and extrapolated by the rational curve $y=y_{0}+\frac{b\cdot x^{d}}{1+c\cdot x}$. The extrapolated value of the spin operator scaling exponent is $2\Delta_{\textmd{S},\infty}=3.9\pm 0.13$, which agrees with the manual envelope linear fitting shown in Fig. 3(a), as well as the analytical estimate $2\Delta_{\textmd{adj}}\in(3.5,3.9)$ mentioned above. Here, the error is obtained from two sources. The first is the errors in the fit curve parameters, which are propagated to the extrapolated value through the differentiation of the curve equation. The second source is the average deviation of the scattered data points from the fit curve. The two sources are considered independent for simplicity.
@@
Fig. 3(b) shows log plot of the bond-bond correlation $C_{B}(r,0)$ as a function of real space distance $r$ at $J=1.25,K=1$, for various lattice sizes. The data of the bond correlation generally have a larger variance than the spin correlation function shown in Fig. 3(a), which is due to the explicit gauge field dependence in the bond operator $B_{i}$. But as the system sizes increases to beyond $L\gtrsim 30$, the decaying behavior of the bond-bond correlation tends to be stable around a straight line, which again indicates a power law decay. The manual linear fitting shown as the dark solid line yields a power of around $3.7$, which again agrees with the scaling dimension of the fermion bilinear operators $2\Delta_{\textmd{adj}}\in(3.5,3.9)$. Similar to the data processing of the spin-spin correlation data described above, the bond-bond correlation data $C_{\textmd{B}}(r,0)$ are also fitted and extrapolated to $1/L\to 0$ as shown in Fig. 3(e), which gives the bond operator scaling exponent of $2\Delta_{\textmd{B},\infty}=3.8\pm 0.2$ and agrees with the analysis above.
@@
As analyzed in the above section and shown in Figs. 3(a) and 3(b), the fact that both spin and bond correlation functions display power law decay with powers consistent with the scaling dimension of the CFT fermion bilinear operators supports the existence of the $U(1)$ DSL phase in non-compact QED3 at $J=1.25$. This also agrees with a recent DQMC study on the same model [11].
@@
Fig. 3(c) shows the log plot of the magnetic flux correlation function $C_{\textmd{flux}}(\tau)$ as a function of imaginary time $\tau$ for various lattice sizes, at $J=1.25,K=1$. It can be seen that, as system size increases, the curves become increasingly linear before entering the noise level at around $2\times 10^{-3}$. Such linear tendency is manually fitted with the dark solid line, which yields a power of around $4.0$. In contrast, the black dashed line shows the $\tau^{-3.0}$ behavior, which originates from the non-conformal flux-flux correlation in a $U(1)$ gauge field theory [11]. This is clearly distinct from the $\tau^{-4.0}$ behavior originating from the conserved current’s scaling dimension $\Delta_{J}=2$, as mentioned above. Thus, the power law decay in the flux correlation function with power $4.0$ also indicates the existence of $U(1)$ DSL phase in $J=1.25$ in the non-compact QED3. Similar to the data processing of the spin-spin correlation data described above, the flux-flux correlation data $C_{\textmd{flux}}(\tau)$ are also fitted and extrapolated to $1/L\to 0$ as shown in Fig. 3(f), which gives the flux operator scaling exponent of $2\Delta_{\textmd{flux},\infty}=3.8\pm 0.7$ and agrees with the analysis above. Here, however, the $1/L\to 0$ extrapolation quality is not as good as that of the fermion bilinears shown in Fig. 3(d, e), and as a result, the error bar of extrapolated exponent is fairly large. This is because the slopes of the $C_{\textmd{flux}}(\tau)$ curves seem to display oscillations at finite lattice sizes, as shown in Fig. 3(c). This indicates that to obtain more accurate flux dynamic correlation $C_{\textmd{flux}}(\tau)$, the simulation might need to be pushed to larger lattice sizes.
@@ II Discussion
Since the HQMC gives the full space-time correlation of the fermion bilinear operators, one immediate future direction is to investigate the spectral information inside the $U(1)$ DSL phase from the stochastic analytic continuation of the HQMC data [11]. As we have both large $N_{\tau}$ and $V_{s}$, the momentum and energy resolution of the obtained spectra (for example the magnetic spectra from spin-spin correlation function) will be much higher than those from the DQMC [11]. In this way, the weight distribution and the momentum dependence of the spinon continuum spectra of $U(1)$ DSL can be revealed, and hopefully connection with experiment results on triangular lattice antiferromagnets YbZn2GaO5 [7, 66] and the $A$-YbSe2 delafossites [55, 56] and kagomé antiferromagnet YCu3(OD)6Br2[Br1-x(OD)x[69, 70, 27, 61, 60, 38, 71] can be made.
Moreover, by adding magnetic field to the QED3 model, one can investigate how the flavor chemical potential will change the $U(1)$ DSL phase and understand the emergent gauge flux in magnetized Dirac spin liquids. For example, recent DQMC work [11] and analytic works [21, 20] have explored the various symmetry-breaking phases out of $U(1)$ DSL under magnetic field. In particular, a novel chiral flux phase, where the system spontaneously generates a net emergent orbital magnetic flux (deviated from the $\pi$-flux at zero field case) has been seen from smaller size DQMC simulation. This spontaneous orbital magnetic flux changes the behavior of the spinons, and make them form the relativistic Landau levels — energy levels normally seen in electrons moving in a magnetic field. Remarkably, the system also exhibits a gapless photon-like mode [11], a signature of the coherent emergent gauge field, which can be potentially detected in inelastic neutron scattering experiments. Our GPU-accelerated HQMC algorithm can be applied to further study such system with rich phase diagram in the presence of magnetic field and gauge field fluctuations at finite temperatures. One possible direction is to investigate the effect of different filling of the spinon Landau levels in the magnetized DSL phase [62], and establish possible connections to experimentally observed unusual magnetic oscillations in the YCu3(OD)6Br2[Br1-x(OD)x] kagomé antiferromagnet [71].
@@ III Methods
As introduced above in the Results section, the model we study consists of a fermionic Lagrangian $L_{F}$, and a gauge field Lagrangian $L_{B}$. And we consider two $U(1)$ gauge field Lagrangians, namely compact $U(1)$ gauge field $L_{B}^{\textmd{c}}$ and non-compact $U(1)$ gauge field $L_{B}^{\textmd{nc}}$ as introduced in the Results section. The discretized version of $\operatorname{curl}\phi$ for a spatial plaquette $\square$[16, 23] , which is also seen as magnetic flux through the plaquette, is defined as $\theta^{xy}_{\square}=\operatorname{curl}\phi=\sum_{ij\in\textmd{edge}(\square)}\phi_{ij}$, where $ij$ is along the positive direction, and $\phi_{ij}=-\phi_{ji}$, as illustrated in the bottom right cube in Fig. 1. Similarly, in the cubic space-time lattice, the curl of a temporal plaquette, i.e. $\theta^{\tau x}$ or $\theta^{\tau y}$, can also be defined. If we fix the gauge field of temporal bonds to be zero, then the curl of a temporal plaquette is $\pm(\phi_{ij,\tau+1}-\phi_{ij,\tau})$, which consists of the $\tau$-derivative term in the gauge field Lagrangians. It is then clear that $L_{B}^{\textmd{c}}$ has $2\pi$-periodicity in the magnetic flux: $\theta\to\theta+2n\pi,n\in\mathbb{Z}$. Thus, $L_{B}^{\textmd{c}}$ describes a compact $U(1)$ gauge field theory in the sense that the magnetic flux is physical only in the range of $(-\pi,\pi]$, while the other values of $\theta$ are unphysical and need to be wound into the $(-\pi,\pi]$ interval. Due to $2\pi$-periodicity (the fermionic Lagrangian $L_{F}$ also has $2\pi$-periodicity by definition [41]), the compact QED3 is known to host magnetic monopoles, which contains distinct physics from a non-compact theory e.g. $L^{\textmd{nc}}_{B}$ in Eq. (2), where the first term breaks the $2\pi$-periodicity. For example, the proliferation of magnetic monopoles in compact QED3 will lead to the emergence of a mass gap in the gauge field, through the dual superconductor mechanism [48, 49, 39], while the non-compact QED3 is generally believed to remain deconfined. We therefore focus on the results on non-compact QED3 in the main text and leave the compact QED3 results in the Supplemental Materials (SM) [2]. Although from the algorithmic point of view, there is no difference in their implementation and performance of the corresponding simulations.
@@
is of size $N_{\tau}V_{s}\times N_{\tau}V_{s}$. Furthermore, it can be shown [67] that $\operatorname{det}M_{\alpha}\in\mathbb{R}$ and $\operatorname{det}M_{\uparrow}=\operatorname{det}M_{\downarrow}$, thus the simulation is free from the sign problem. Based on the formula above, the partition function can then be expressed as
@@
In this subsection, we will briefly introduce the HQMC algorithm applied to simulate the partition function Eq. (9). This algorithm originates from lattice quantum chromodynamics simulation [19], and later finds its application in a broader class of statistical problems [10]. An application in the Hubbard model setting of condensed matter physics can be found in Ref. [8]. In the most recent progress, the method is significantly accelerated with the help of GPU in the study of spin-fermion models and the associated non-Fermi-liquid and quantum critical metallic states [47, 42].
@@
Here, the Gaussian integral of the complex fields $\eta\in\mathbb{C}^{V_{s}N_{\tau}}$ has been conversely used, and the identity $\operatorname{det}M_{\uparrow}=\operatorname{det}M_{\downarrow}\in\mathbb{R}$ is assumed, and proved in Ref. [67]. Thus, the fermionic flavor $\alpha$ is omitted.
@@
First, the Hamiltonian dynamics has the property that the phase space volume is conserved, according to the Liouville theorem [8], which ensures that $H(\phi_{0},p_{0},\eta)=H\left(\phi_{t},p_{t},\eta\right)$ and that the proposal variables $(\phi_{t},p_{t})$ are almost always accepted, if the differential equation is solved exactly.
Second, we apply the well-known Leapfrog method [8] to solve the differential equation Eq. (14). It has the property of time reversibility, which, together with the Metropolis-Hasting rule, guarantees that the Monte Carlo updates fulfill the detailed balance condition [8]. The Leapfrog method is described as follows. At the $n$-th step, to evolve a given $(\phi_{n-1},p_{n-1})$ for a time step $\Delta t$,
@@
Third, since the update scheme in HQMC is based on solving the Hamiltonian differential equation, it is naturally continuous. As a result, the random walk in HQMC is limited to a continuous region the configuration space. But in the QED3 model studied in this paper, there are two non-continuous regions, namely those with $(\det M_{\uparrow},\det M_{\downarrow})=(+,+)$ and $(\det M_{\uparrow},\det M_{\downarrow})=(-,-)$. They are separated by the gauge configurations where $\det M\approx 0$, which themselves have diminishing contribution in the partition function Eq. (9), but prevents the HQMC update from going from one region to the other. This leads to the ergodicity issue, shared by the application of HQMC on other models [8] as well. One common solution is to introduce another legitimate update that is ergodic. Then the combination of the multiple updates will cure the ergodicity issue in HQMC [10]. In Ref. [8], the solution of this issue is to decouple the Hamiltonian into two channels and then to complexify the configuration space, which is more problem-specific. In the QED3 studied here, we find that in the small $J$ region, which is the $U(1)$ DSL phase region and the focus of this paper, the ergodicity issue is numerically negligible. So we directly apply the HQMC algorithm in this region. The validity is benchmarked by the DQMC algorithm and is shown in Sec. II of SM [2].
@@
For a given bosonic configuration $\phi$, the fermionic observables are easily computed by using the standard stochastic estimation procedure [19, 14]. More specifically, the single-particle Green’s function, $G_{ij}=\langle c_{i}(\tau)c^{\dagger}_{j}(0)\rangle$, which is obtained from the inversion of the $M$ matrix, $M^{-1}$. The direct inversion of $M$ is numerically expensive and scales as $O((V_{s}N_{\tau})^{3})$. Alternatively, the standard stochastic procedure to compute $M^{-1}$, as an iterative method, is usually more efficient. In this procedure, first sample a random vector $\xi\in\mathbb{C}^{V_{s}N_{\tau}}$, whose entries are i.i.d. and have the orthogonality property $\mathbb{E}_{\xi}\xi^{*}_{i}\xi_{j}=\delta_{ij}$. Then, $\sum_{l}\left[M^{-1}\right]_{il}\xi_{l}\xi^{*}_{j}$ will give an unbiased estimation of $\left[M^{-1}\right]_{ij}$, since
@@
Here, the random vector can be drawn from any distribution that satisfies $\mathbb{E}_{\xi}\xi^{*}_{i}\xi_{j}=\delta_{ij}$, including Gaussian distribution, Rademacher distribution, etc [14]. In practice, we compute $\sum_{l}\left[\left(M^{\dagger}M\right)^{-1}\right]_{il}\left[M^{\dagger}\xi\right]_{l}\xi^{*}_{j}$, instead, where $R:=M^{\dagger}\xi$ is numerically cheap to compute and $\left[M^{\dagger}M\right]^{-1}R$ can be computed generically using a pCG algorithm. In this way, the Green’s function can be obtained, and other more complicated correlation functions can be further obtained from Green’s function via Wick decomposition.
@@
has a time complexity of $O(N^{2})$, where $N=V_{s}N_{\tau}$. Here, we follow the Ref. [14], and apply the fast Fourier transformation (FFT) instead of directly eevaluating this average.
@@
where $\mathcal{F}$ is FFT operation. Then the time complexity is reduced to $O(N\log N)$, with $N=V_{s}N_{\tau}$. Moreover, as introduced in Ref. [14], each random vector among $N_{\textmd{rv}}$ random vectors can serve as either $\xi$ or $\xi^{\prime}$ in the example above. Thus, totally $f(N_{\textmd{rv}})=\binom{N_{\textmd{rv}}}{2}$ random vectors can be effectively used in the stochastic estimation. In our benchmark, $N_{\textmd{rv}}=40$ is sufficient to reduce the error of fermionic bilinear’s correlation function to a moderate level.
@@
As introduced above, the key step in reducing the cubic complexity of computing the determinant $\det M$ or the inversion $M^{-1}$ is to convert it to the evaluation of $\left[M^{\dagger}M\right]^{-1}v$, which is equivalent to solving a linear system $OX=v$. Here, $O=M^{\dagger}M$ is a Hermitian positive semi-definite matrix; thus, the equation can be efficiently solved by the iterative preconditioned conjugate gradient (pCG) method [8].